# Exciting Contact Modes in Differentiable Simulations for Robot Learning

### v1

@@ I Introduction
In order to search for rich contact data for effective parameter learning, we require the identification of key contact interactions and a way to plan control inputs to explore for contacts. To achieve this, we propose an experimental design approach [7, 8, 9] that is based on contact-implicit optimization that optimizes contact modes which maximizes a contact-aware Fisher information [10, 11] which directly reduces a lower-bound on parameter uncertainty [12, 9]. Our approach integrates contact planning [13, 14, 15, 5] with the contact-aware information metric to facilitate smoothness in the computation of gradients that are useful for down-stream learning tasks. We show that our approach is able to search for contact modes that maximizes information which improves learning of unknown parameters of interest, thus guiding and enhancing the robot learning process.
@@ III Results and Discussion
IV Conclusion
We present an experimental learning approach that identifies and searches for information-rich contacts in differentiable simulations by actively planning and exciting contact modes to improve parameter learning. Compared to random sampling methods, we show improved parameter learning and richer information from contact modes with respect to Fisher information parameters.

### v2

@@ I Introduction
In order to search for rich contact data for effective parameter learning, we require the identification of key contact interactions and a way to plan control inputs to explore for contacts. To achieve this, we propose an experimental design approach [7, 8, 9] that is based on contact-implicit optimization that optimizes contact modes which maximizes a contact-aware Fisher information [10, 11] which directly reduces a lower-bound on parameter uncertainty [12, 9]. Our approach integrates contact planning [13, 14, 15, 16, 17, 18, 5] with the contact-aware information metric to facilitate smoothness in the computation of gradients that are useful for down-stream learning tasks. We show that our approach is able to search for contact modes that maximizes information which improves learning of unknown parameters of interest, thus guiding and enhancing the robot learning process.